A 58-kg ice skater moves across the ice at a constant speed of 2.0 m/s. She is caught by her 74-kg partner, and then the pair continues to glide together. He is at rest when he catches her, and immediately afterward they both coast.A skater follows the girl at a constant speed of 1.0 m/s. What is the magnitude of the girl's change in momentum in the reference frame of the skater?
1
Expert's answer
2020-07-20T15:02:22-0400

According to the law of conservation of momentum the final impulse of the girl is "2\times58 - 74 = 42" , which means its speed is "\frac{42}{58} = 0.72" m/s. Therefore, in the skater's reference system, it's initial speed is 2 m/s, and the final speed is 0.72 - 1 = - 0.28 m/s, so the change in its momentum will be "58\times(2 + 0.28) = 132.2" kgm /s.